OpenStudy (anonymous):

You invested a total of $6,000 in two funds earning 7% and 9$ simple interest. During 1 year, the two funds earned a total of $500 in interest. How much did you invest in each fund?

OpenStudy (anonymous):

Not quite sure what I would do first

OpenStudy (anonymous):

Need to set up 2 equations and then use elimination or substitution

OpenStudy (anonymous):

Any certain formulas?

OpenStudy (anonymous):

No..what would be your first equation using A for one fund and b for the other fund?

OpenStudy (anonymous):

Well we know it has to equal 6,000

OpenStudy (anonymous):

Good so A + B = 6000 would be your first equation how about the second equation?

OpenStudy (anonymous):

Hint: let A represent the 7% fund and B represent the 9% fund...

OpenStudy (anonymous):

Ok so the first equation would be .07+.09=6,000

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

A + B =6000 that is the first equation

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

So then would be be

OpenStudy (anonymous):

Wait no nvm what I was about to do was wrong

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

But to start its .07+.09= something right?

OpenStudy (anonymous):

Very Good...So .07 of what and .09 of what???

OpenStudy (anonymous):

Not quite sure

OpenStudy (anonymous):

(Sorry math isn't my strongest subject)

OpenStudy (anonymous):

OK how do you determine 3% interest of say 1000 dollars?

OpenStudy (anonymous):

Multiply?

OpenStudy (anonymous):

Yes .03 times 1000 = 30 which would be the amount of interest earned. So looking at the question .07(something) + .09 (something) = Interest Earned...Can you see? Now what would be the somethings we need to fill in?

OpenStudy (anonymous):

So lets say something is A right?

OpenStudy (anonymous):

So now it has to be what does it equal

OpenStudy (anonymous):

One something is A and the other something would be B..because that represents the two funds...so you have .07(A) +.09(B) = What?

OpenStudy (anonymous):

6,000?

OpenStudy (anonymous):

No A + B =6000 which is the principal amount you started with... .07(A) + .09(B) is the interest earned....how much total interest did you earn?

OpenStudy (anonymous):

Oooooooh 500

OpenStudy (anonymous):

So now you have 2 equations which look like what?

OpenStudy (anonymous):

A way to get the answer?

OpenStudy (anonymous):

Yes show me the two equations

OpenStudy (anonymous):

Ok so the first one is

OpenStudy (anonymous):

A+B=6,000

OpenStudy (anonymous):

2nd one is

OpenStudy (anonymous):

.07A+.09B=500

OpenStudy (anonymous):

Great..do you know how to solve by substitution?

OpenStudy (anonymous):

Ummm

OpenStudy (anonymous):

I don't get what you mean?

OpenStudy (anonymous):

what we are going to try to do is create an equation with only one variable (either A or B) so we can solve. Let's try to solve for B...

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

In you rfirst equation A+ B = 6000 how can we show it as an equation where B= (something)

OpenStudy (anonymous):

Move A to the other side right?

OpenStudy (anonymous):

Correct..subtract A form both sides and what do you have?

OpenStudy (anonymous):

So A-A+B=6,000-A?

OpenStudy (anonymous):

B=6,000-A

OpenStudy (anonymous):

Great now we have B=6000- A....now look at your second equation of .07(A) + .09(B) = 500..we are going to solve by substitution so what can we substitute based on our first equation

OpenStudy (anonymous):

I made an error earlier we are going to solve for A...

OpenStudy (anonymous):

Oh then just the opposite

OpenStudy (anonymous):

A=6,000-B

OpenStudy (anonymous):

No

OpenStudy (anonymous):

We have the equation .07(A) +.09(B) = 500 and we know B= 5000-A..we simply substitute the value we have determined for (B) into every value for B in the equation .07(A) + .09(B) =500....

OpenStudy (anonymous):

Oh wait we had it right at first?

OpenStudy (anonymous):

So you end up with .07(A) + .09(5000-A) = 500...do you see where I substituted?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now distribute

OpenStudy (anonymous):

Very good and solve for A

OpenStudy (anonymous):

Kk give me a minute to solve it

OpenStudy (anonymous):

Ok so A=2,000

OpenStudy (anonymous):

So now we substitute that into A+B=6,000

OpenStudy (anonymous):

To solve for B?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So the answer for B is 4,000?

OpenStudy (anonymous):

Very good...you can always verify your answers by plugging them into both equations and making sure they work...Great job

OpenStudy (anonymous):

Ok and 1 more quick thing

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

When we did a problem in class the answer came out to 18 quarters and my teacher said I had to put it into dimes as well someone immediately said 32 dimes and he said that is correct then class ended so I didn't know how he got that... How exactly did he get 32?

OpenStudy (anonymous):

Unless I saw the problem I could not tell you because 18 quarters do not equal 32 dimes. I am sure this is another substitution problem as coins are often used. Do you have the problem handy?

OpenStudy (anonymous):

Yeah I got it right here I'll type it

OpenStudy (anonymous):

A person has 50 coins in dimes and quarters with a combined vaule of 7.70. Determine the number of coins of each type.

OpenStudy (anonymous):

So waht are your two equations going to look like....let D= dimes and Q = quarters.

OpenStudy (anonymous):

Well what I did was value+value=total value because we had a similar problem that we had to do the that

OpenStudy (anonymous):

So I put down 25x + 10(50-X)= 770

OpenStudy (anonymous):

Wait I am stupid

OpenStudy (anonymous):

Nvm I figured it out

OpenStudy (anonymous):

No you are on the right track..

OpenStudy (anonymous):

I never thought of this lol 18+32=50....

OpenStudy (anonymous):

That's how he got 32....

OpenStudy (anonymous):

Well X=18

OpenStudy (anonymous):

I feel really really really dumb

OpenStudy (anonymous):

Never feel like that you do well....sometimes it takes a second to see it

OpenStudy (anonymous):

Yeah just to make sure I did this right

OpenStudy (anonymous):

It's the same kind of problem I asked at first

OpenStudy (anonymous):

I just wanna make sure I have it down

OpenStudy (anonymous):

You invested a total of $30,000 in two funds earning 8.5% and 10% simple interest. During 1 year, the two funds earned a total of $2,700 in interest. How much did you invest in each fund?

OpenStudy (anonymous):

So same concept

OpenStudy (anonymous):

A+B=30,000

OpenStudy (anonymous):

Exact same

OpenStudy (anonymous):

Does it really matter which I solve for at first?

OpenStudy (anonymous):

Say if I want to solve for A this time does that really matter?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

So I can do A or B

OpenStudy (anonymous):

Yes but what will your second equation look like?

OpenStudy (anonymous):

.085A+.10B=2,700

OpenStudy (anonymous):

Correct..

OpenStudy (anonymous):

Ok! Thank you so much!

OpenStudy (anonymous):

If I have another problem think you could help? But the rest of the problems are pretty simple...

OpenStudy (anonymous):

Always glad to help...on the coin question I just wanted to make sure that you have that set up correct...It is based on the number of coins so your two equations should look like this.... D + Q = 50 and .10(D) +.25 (Q) = $7.70 where D and Q represent the number of each coin...

OpenStudy (anonymous):

What is the other problem